### v1

# Talk2PC: Enhancing 3D Visual Grounding through LiDAR and Radar Point Clouds Fusion for Autonomous Driving

Abstract
Embodied outdoor scene understanding forms the foundation for autonomous agents to perceive, analyze, and react to dynamic driving environments. However, existing 3D understanding is predominantly based on 2D Vision-Language Models (VLMs), collecting and processing limited scene-aware contexts. Instead, compared to the 2D planar visual information, point cloud sensors like LiDAR offer rich depth information and fine-grained 3D representations of objects. Meanwhile, the emerging 4D millimeter-wave (mmWave) radar is capable of detecting the motion trend, velocity, and reflection intensity of each object. Therefore, the integration of these two modalities provides more flexible querying conditions for natural language, enabling more accurate 3D visual grounding. To this end, in this paper, we exploratively propose a novel method called TPCNet, the first outdoor 3D visual grounding model upon the paradigm of prompt-guided point cloud sensor combination, including both LiDAR and radar contexts. To adaptively balance the features of these two sensors required by the prompt, we have designed a multi-fusion paradigm called Two-Stage Heterogeneous Modal Adaptive Fusion. Specifically, this paradigm initially employs Bidirectional Agent Cross-Attention (BACA), which feeds dual-sensor features, characterized by global receptive fields, to the text features for querying. Additionally, we have designed a Dynamic Gated Graph Fusion (DGGF) module to locate the regions of interest identified by the queries. To further enhance accuracy, we innovatively devise an C3D-RECHead, based on the nearest object edge. Our experiments have demonstrated that our TPCNet, along with its individual modules, achieves the state-of-the-art performance on both the Talk2Radar and Talk2Car datasets.
Index Terms:
3D visual grounding, vision-language model, 4D radar, LiDAR, multi-sensor fusion, autonomous driving
Figure 1: Overview of Talk2PC pipeline, where the textual prompt guides LiDAR and 4D mmWave radar (radar) to localize the referred object(s).
I Introduction
With the rapid advancement of Autonomous Driving (AD) and Intelligent Transportation Systems (ITS), 3D perception based on multi-sensor fusion has been widely adopted in Autonomous Vehicles (AVs), robotics, and roadside perception systems [1, 2, 3]. As a primary sensor for 3D perception, LiDAR provides accurate positioning and detailed 3D representations of objects, and has been proved suitable for numerous perception tasks in AD [4, 5, 6, 7, 8, 9]. However, LiDAR cannot perceive crucial information such as the motion trend and velocity, which are essential for surrounding entities. Luckily, the mmWave radar (radar) is capable of sensing the distance and orientation of each object, with a detection range that exceeds that of LiDAR [10]. Moreover, it can capture motion and velocity, and operates reliably under adverse weather conditions [11]. The latest 4D radar offers richer point cloud data when compared with conventional [12], showing significant potential for 3D perception as well [13, 14, 15, 16, 17, 18, 19]. Consequently, a considerable amount of research has focused on the fusion of LiDAR and radar for enhanced 3D perception [20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
Besides, as Vision-Language Models (VLMs) are progressively applied to interactive perception and embodied intelligence, they enable AVs and robots to not only perceive scenes but also understand human intentions and locate corresponding objects [30, 31]. Nonetheless, the current advancements have predominantly focused on the integration with visual modalities [32], while there has been limited interaction with 3D point cloud sensors, particularly in the domain of language-guided multi-sensor fusion for 3D visual grounding in traffic scenarios. For instance, Talk2Car [33] proposes a benchmark of 2D visual grounding on the image plane from the view of a driving car. WaterVG [34] and NanoMVG [35] focus on the 2D visual grounding based on camera-radar fusion. Cheng et al. [36] extended Talk2Car to LiDAR-based 3D visual grounding while Talk2Radar [37] explores the capacity of 4D radars on 3D visual grounding. Based on the aforementioned, we empirically argue that 2D image plane is inadequate for accurately perceiving the precise spatial coordinates of objects, particularly in terms of depth information. Moreover, the object attributes captured by a single sensor are inherently limited. When the content described by the prompt does not correspond to the information perceived by the sensor, it will lead to incorrect model optimization and parameter learning.
To address the aforementioned issues, in this paper, we propose an exploratory approach that leverages natural language guidance for LiDAR and radar fusion in a dual-sensor framework for 3D visual grounding, as illustrated briefly in Fig. 1. Given that the description patterns of object attributes in prompts are not fixed, it necessitates the dynamic weighting for the importance of the LiDAR and radar in relation to the prompt. Unlike the previous methods that employed static fusion strategies based on convolutions for LiDAR and radar [38, 26, 23, 27, 39], we design a novel Bi-directional Agent Cross Attention (BACA). This approach performs multi-scale feature fusion in the encoder stage, adopting LiDAR and radar as the primary information source for the query alternately. Notably, BACA significantly alleviates the computational complexity associated with vanilla cross-attention while enabling efficient dynamic modeling. Furthermore, to mitigate the issue of false positives from point cloud sensors in 3D visual grounding and to correlate multiple point cloud objects mentioned in the prompt, we propose a novel text-to-point-cloud fusion strategy, termed Dynamic Gated Graph Fusion (DGGF). The limitation of previous Gated Graph Fusion (GGF) [37] is that it uses a static graph construction method, all features construct the same graph, and do not consider the similarity of nodes, which makes it impossible to adjust the graph structure according to the image content, limiting the advantages of graph-based models. In contrast, DGGF constructs a linguistically-conditional candidate region modeling process by utilizing a dynamic axial graph neural network based on feature gating. Furthermore, to prevent prediction errors caused by the absence of object center point clouds and to align the depth information as closely as possible with the descriptions in the prompt, we propose a 3D visual grounding prediction head, termed Corner3D-RECHead (C3D-RECHead). This head uses the region with denser point clouds near the corner of the object closest to the ego vehicle as the anchor centroid.
The main contributions of this paper are summarized as below:
  • We propose a novel model, named TPCNet, which performs 3D visual grounding by leveraging textual semantics to guide dual point cloud sensors. TPCNet is capable of adaptively and dynamically aligning and fusing linguistic features with heterogeneous point cloud data. It achieves state-of-the-art performance on both the Talk2Radar and Talk2Car datasets.
  • We propose the Bi-directional Agent Cross Attention (BACA) to fuse the features of LiDAR and radar with sufficient consideration of respective physical characteristics, which enables TPCNet to embed radar representations of object motion alongside LiDAR’s refined 3D geometric understanding of the environment.
  • We propose Dynamic Gated Graph Fusion (DGGF), which constructs a dynamic axial graph for the adaptive fusion of point cloud features in the lingusitic feature space, while filtering regions of interest.
  • In alignment with the physical properties of point cloud sensors, we propose a 3D visual grounding head called C3D-RECHead, centered on the closest edge between the object and the ego-vehicle. C3D-RECHead effectively reduces the object miss detection rate while significantly enhancing localization precision in samples where the prompt includes numeric depth information.
Figure 2: The architecture of our proposed TPCNet. In TPCNet, Bidirectional Agent Cross Attention (BACA), Dynamic Gated Graph Fusion (DGGF) and C3D-RECHead are three modules proposed by us, which are illustrated in detail in Sub-section III-C, III-D and III-E.
II Related Works
II-A 3D Object Detection with LiDAR-Radar Fusion
Leveraging the high-precision 3D sensing capabilities of LiDAR and the all-weather robustness and long-range sensing strengths of radar, the fusion of LiDAR and radar has been regarded as a complementary solution for 3D object detection. The works in [26, 23, 40, 38] propose the bi-directional radar-LiDAR fusion module, which is based on the convolution-based attention with the product of feature maps to weigh the respective features dynamically. Besides, in [26], the Interaction-Based Multi-Modal Fusion (IMMF) module is proposed for radar-LiDAR fusion based on bi-directional convolutional weighting. Meng et al. [25] introduced cross fusion to generate the pseudo radar feature upon the guidance of LiDAR. Wang et al. [27] proposed a two-stage LiDAR-radar fusion method to obtain the high-quality composite point cloud features. Building on the above, we observe that, first, several approaches adopt LiDAR-guided radar strategies to complement the features of the two modalities. Second, some methods employ bi-directional fusion to enhance missing features from both modalities while highlighting the object regions. However, these methods are primarily static convolution-based, which demonstrates limited generalization to sensor variations across different scenarios. To address this, the works such as [20, 28] propose attention-based fusion methods to enable dynamic cross-attention and information exchange between modalities. Nonetheless, these attention-based approaches typically operate on high-dimensional features, causing a quadratic increase in computational complexity and significant model parameter overhead, making their representational efficiency compromised. Here, we propose the Bi-directional Agent Cross Attention with linear complexity to take LiDAR and radar as the double information sources of 3D geometric context and motion features to obtain the abundant features that can be aligned with textual prompts.
II-B Outdoor Visual Grounding
Visual grounding, a vision-language task aiming to localize objects based on natural language prompts [41], has gained significant attention in outdoor applications ranging from autonomous vehicles to embodied intelligence [42]. Recent advances extend beyond traditional camera-based 2D grounding to multi-sensor and 3D spatial understanding [43].
In 2D domain, Guan et al. [35] developed a lightweight vision-radar fusion model for waterway monitoring in USVs, while their later work [34] established a multi-task benchmark with two-stage fusion of camera, radar, and language features. However, these approaches remain limited to 2D planes and lack comprehensive 3D spatial reasoning capabilities.
The field has recently evolved toward 3D outdoor grounding, driven by autonomous navigation demands. Cheng et al. [36] pioneered LiDAR-based 3D grounding for driving scenarios, whereas Wu et al. [44] achieved multi-camera object tracking through language prompts and Zhan et al. [45] proposed a monocular-based 3D REC baseline model called Mono3DVG-TR. Radar-based solutions have also emerged, with Guan et al. [37] establishing a 4D radar benchmark on the VoD dataset [46]. Despite these advancements, the current research exhibits three limitations: (1) insufficient integration of complementary sensors, (2) limited adaptability to dynamic textual guidance, and (3) under-explored human-robot interaction paradigms for UGVs.
Our work bridges these gaps through two key contributions. First, we introduce a unified 3D grounding framework to support pluggable point cloud sensors (e.g., LiDAR and radar) with modality-adaptive fusion. Second, we develop dynamic attention mechanisms that automatically adjust sensor weighting based on textual semantics, outperforming conventional sequential fusion approaches. The proposed Talk2PC system demonstrates superior performance in both autonomous driving scenarios and interactive robot navigation tasks.
III Methodology
III-A The Overall Pipeline
Fig. 2 presents the overall architecture of our proposed TPCNet. TPCNet receives inputs from two perception sensors: LiDAR and 4D mmWave radar, both represented in the form of point clouds. Each LiDAR point cloud contains 3D spatial location information $(x,y,z)$ and intensity $(ins)$. Similarly, each 4D mmWave radar point cloud includes 3D spatial location information $(x,y,z)$ aligned to the LiDAR coordinate system, along with radar cross-section $(rcs)$ and compensated radial velocity $(v)$. Additionally, TPCNet accepts textual prompts from the agent to guide the localization of objects in the scene using both sensors.
For the LiDAR and radar inputs, each passes through a pillar-based backbone network, generating multi-scale features at three levels: LiDAR-based features $f^{l}_{i,i\in\{1,2,3\}}$ and radar-based features $f^{r}_{i,i\in\{1,2,3\}}$. The natural language instructions provided by the agent are processed through a text encoder to extract textual features.
The multi-scale LiDAR and radar features are then input into the Bidirectional Agent Cross Attention (BACA) module, where BACA is performed to produce three levels of LiDAR-Radar fusion features $lr_{i,i\in\{1,2,3\}}$. These fused features are further integrated with the textual features in the Dynamic Gated Graph Fusion (DGGF) module, yielding language-embedded contextual features $lc_{i,i\in\{1,2,3\}}$ at three scales.
Subsequently, these features are passed through an FPN to enhance multi-scale feature representation, resulting in the refined features $f_{i,i\in\{1,2,3\}}$. Finally, these features are fed into the C3D-RECHead, which localizes the queried object within the point cloud scene based on the textual instruction, outputting the object’s 3D position, size, and orientation.
III-B Backbones of LiDAR, Radar and Textual Instruction
During the feature encoding stage, the LiDAR and radar point clouds are first processed through pillar-based backbones [47], encoding the point clouds into Bird’s-Eye-View (BEV) pillar features $f^{l}$ and $f^{r}$. Each backbone produces features across three stages, where the feature shape for a single stage is given by $\{f^{l},f^{r}\}\in\mathbb{R}^{C\times H\times W}$.
For the textual instructions provided by the agent, we utilize the text encoder in PointCLIP [48] for encoding instead of standalone text encoders (e.g., BERT [49] or RoBERTa [50]), as these lack an understanding of point cloud information. PointCLIP leverages the powerful contrastive learning capabilities of CLIP [51] to embed point clouds and text into a shared representation space, enabling robust cross-modal representation learning. Using PointCLIP, we obtain the textual features $pl\in\mathbb{R}^{C\times L}$.
III-C Bidirectional Agent Cross Attention
Figure 3: Demonstration of the importance of BACA for 3D visual grounding based on LiDAR and radar.
LiDAR provides high-precision three-dimensional spatial information, while radar excels in diverse weather conditions and is highly effective at detecting the motion of objects. By fusing the data from these two sensors, a more comprehensive environmental understanding can be achieved. Moreover, the textual instructions used to describe the object do not necessarily encapsulate all relevant aspects, such as the semantic meaning, velocity, direction, and depth of the object simultaneously. As such, it is crucial for both LiDAR and radar to dynamically extract features to ensure that the correct attributes are aligned with the textual instructions. Furthermore, for a dual-sensor system, maximizing the utilization of contextual features captured by each sensor and exchanging this information serves to mitigate the information asymmetry inherent in one-way fusion. Additionally, bidirectional fusion facilitates mutual verification between the sensors. Based on these considerations, we propose an efficient and rapid bidirectional fusion module: Bidirectional Agent Cross Attention (BACA).
Figure 4: The detailed structure of Bidirectional Agent Cross Attention for the fusion of LiDAR and radar feature. Here is the instance for one-stage features.
As depicted in Fig. 4, the features from LiDAR $f^{l}\in\mathbb{R}^{C\times H\times W}$ and Radar $f^{r}\in\mathbb{R}^{C\times H\times W}$ are initially augmented with positional encoding $\mathtt{PE}$ [52]. These enhanced features are subsequently passed through three linear feedforward modules to generate triplet features. In this context, $Q_{L}$, $K_{L}$, and $V_{L}$ represent the query, key, and value matrices for the LiDAR, respectively, while $Q_{R}$, $K_{R}$, and $V_{R}$ denote the corresponding matrices for the radar. The initialization process described above is detailed in Equations (1) to (6) as below:
$Q_{L}=(f^{l}+\mathtt{PE})\mathbf{W}_{QL},Q_{L}\in\mathbb{R}^{C\times H\times W},$ (1)
$K_{L}=\mathtt{Flat}((f^{l}+\mathtt{PE})\mathbf{W}_{KL}),K_{L}\in\mathbb{R}^{C\times L\ (L=H\times W)},$ (2)
$V_{L}=\mathtt{Flat}((f^{l}+\mathtt{PE})\mathbf{W}_{VL}),V_{L}\in\mathbb{R}^{C\times L\ (L=H\times W)},$ (3)
$Q_{R}=(f^{r}+\mathtt{PE})\mathbf{W}_{QR},Q_{R}\in\mathbb{R}^{C\times H\times W},$ (4)
$K_{R}=\mathtt{Flat}((f^{r}+\mathtt{PE})\mathbf{W}_{KR}),K_{R}\in\mathbb{R}^{C\times L\ (L=H\times W)},$ (5)
$V_{R}=\mathtt{Flat}((f^{r}+\mathtt{PE})\mathbf{W}_{VR}),V_{R}\in\mathbb{R}^{C\times L\ (L=H\times W)},$ (6)
where $C$ is the number of channel while $H$ and $W$ denotes the height and width of the feature map. $L$ is the product result of $H$ and $W$. $\mathtt{Flat}(\cdot)$ denotes the flatting operation along the spatial dimension for the image-like feature map.
Furthermore, we generate the context agent features of LiDAR and radar, $A_{L}\in\mathbb{R}^{C\times h\times w}$ and $A_{R}\in\mathbb{R}^{C\times h\times w}$, by exerting the adaptive pooling to the two query matrices $Q_{L}$ and $Q_{R}$, where two context agent features $A_{L}$ and $A_{R}$ maintain the basic contextual structures while reducing the computational cost for the downstream calculation on the softmax attention. Based on the above, we leverage softmax attention $\mathtt{Attn}(\cdot)$ to fuse the features of two sensors. Exactly, for the generation process of LiDAR-driven geometric feature $f_{lg}$, the lidar context agent feature $A_{L}$ firstly serves as the agent of $Q_{L}$, which aggregates the global contextual information in $K_{R}$ and $V_{R}$ provided by radar. Based on the above, we obtain the LiDAR-Agent Cross Feature $f_{lc}$. Subsequently, we take $A_{L}$ as the key while $f_{lc}$ as the value in the second softmax attention with the orginal $Q_{L}$ as the query. Here, we broadcast the global contextual information upon the agent feature $A_{L}$ to every query tokens, where such process avoids the direct calculation of pairwise similarities between query and key features, maintaining the information exchange via agent feature. Likewise, the generation process of Radar-driven motion feature $f_{rm}$ follows the same process with LiDAR-driven geometric feature $f_{lg}$. The whole process is shown in Eq. (7) and (8).
[…]
where $\mathtt{AdaPool}(\cdot)$ represents the adaptive pooling operation. $l$ is the product result of $h$ and $w$, where $h$ and $w$ denote the height and width of the agent feature. The values of $h$ and $w$ are both smalled than $H$ and $W$. In this context, the size of the agent feature $l$ is a hyperparameter, significantly smaller than the original LiDAR and Radar feature dimensions. It achieves a linear computational complexity of $O(LlC)$. The complexity of our proposed BACA is much smaller than vanilla cross attention’s $O(L^{2}C)$ while still preserving the global cross-modal fusion capability.
III-D Dynamic Gated Graph Fusion
Figure 5: The construction of the dynamic graph in DGGF. (a) The graph construction in GGF for the green patch of an 8 $\times$ 8 feature map. All red regions will be linked to the green region, irrespective of their similarity. (b) Dynamic graph construction applied to the green region of an 8 $\times$ 8 feature map, which adaptively builds a graph along the axes by utilizing a mask (represented by the blue regions), and ensures that only patches with similar Euclidean distances are connected. The red patches remain disconnected from the green patch, as they do not fall within the scope of the mask.
To efficiently fuse the features of point clouds and textual prompts, as shown in Fig. 2, we employ a cross-modal gating mechanism to integrate the language and point cloud features, and construct dynamic graph aggregation to capture the regions of interest within the point cloud context.
Let the linguistic feature provided by PointCLIP be denoted as $f_{t}$. We first apply the Max-Pooling operation to obtain the compressed linguistic feature $f_{t}^{s}$, which then passes through a feedforward layer with a sigmoid activation function to compute the gating weight $\mathbf{W}_{G}$. In parallel, the LiDAR-Radar fusion feature $lr$ is first augmented with Conditional Position Encoding ($\mathtt{CPE}$), and the resulting position-aware feature is element-wise multiplied by the gating weight $\mathbf{W}_{G}$, yielding the language-conditioned point cloud feature $f_{t}^{lc}$. The detailed process is shown in below:
$f_{t}^{s}=\mathtt{MaxPool}(f_{t}),$ (9)
$\mathbf{W}_{G}=\rho(\mathbf{W}\cdot f_{t}^{s}),$ (10)
$f_{t}^{lc}=\mathbf{W}_{G}\cdot(lr+\mathtt{CPE}(lr)),$ (11)
where $\rho(\cdot)$ denotes the Sigmoid function.
Secondly, to capture the regions of interest within the point cloud encoded by the textual prompt, as well as the surrounding contextual information, we leverage the power of graph structures, which excel at handling data with complex spatial or structural relationships. Given that point cloud data lacks a fixed grid structure, we model the features as a set of nodes within a graph. This approach enhances the ability to model contextual information, thereby improving the understanding of both the spatial distribution and the semantic information of the object. As Fig. 5 shows, compared with the static graph constructed by Sparse Vision Graph Attention (SVGA) [53], which cannot be dynamically adjusted according to the characteristics of different features, we construct the dynamic graph which leverages the axial structure of SVGA to maintain its efficiency while dynamically constructing a more representative graph, changing across the features. To achieve this, we estimate the mean ($\mu$) and standard deviation ($\sigma$) of the Euclidean distances between nodes by utilizing a node subset. This subset is obtained by partitioning the pseudo-image feature map of the point cloud into quadrants and comparing the diagonal pairs of these quadrants. This approach enables the computation of the estimated $\mu$ and $\sigma$ between the feature maps of the original pseudo-image and its diagonally flipped counterpart. The purpose of this estimation is to reduce the number of pairwise comparisons required to obtain $\mu$ and $\sigma$. We avoid directly computing the exact values of $\mu$ and $\sigma$ because doing so would necessitate calculating the Euclidean distance between every individual node and all other nodes in the image, which is computationally expensive. Subsequently, we adopt the row and column-wise connections as in SVGA to further reduce computational complexity. As demonstrated by MobileViG [53], every node should not be considered. If the Euclidean distance between two nodes is less than the difference between the estimated $\mu$ and $\sigma$, we connect the two nodes. DGGF also supports a variable number of connections across different point cloud feature maps, which contrasts with the standard approach in graph neural networks where a fixed number of K Nearest Neighbors (KNN) is used for all images. This is because, within different feature maps, the Euclidean distance between nodes is often smaller than the estimated difference between $\mu$ and $\sigma$. The intuition behind using $\mu$ and $\sigma$ is that nodes within the range of $\mu$$\pm$ $\sigma$ are close to each other and should share information. These values are then used to establish connections between the generated nodes, as shown in Figure 2b. In Fig. 5, we can observe that GGF connects the Region of Interest (RoI) of the object to non-object regions in the image, whereas our proposed DGGF only connects the object region to other object regions in the point cloud feature map.
Algorithm 1 Algorithm of Dynamic Gragh Convolution ($\mathtt{DynConv}$) in DGGF
Input:
$K$: the step length of the connection between nodes;
X: the input feature map with the shape $H\times W$;
Xquadrants: the quadrants of the input flipped across the diagonals;
$m$: the distance of each roll.
1:$\textit{X}_{norm}$leftarrow $\mathtt{norm}$(X, Xquadrants)   $\triangleright$ matrix norm of tensors
2:$\mu$leftarrow $\mathtt{mean}$($\textit{X}_{norm}$)
3:$\sigma$leftarrow $\mathtt{std}$($\textit{X}_{norm}$)
4:while $mK$ $<$ H do
5:  Xrolled $\leftarrow$ $\mathtt{roll}_{down}$(X, $mK$)
6:  dist $\leftarrow$ $\mathtt{norm}$(X, Xrolled)   $\triangleright$ get distance value
7:  if dist $<$ ($\mu$- $\sigma$) then   $\triangleright$ generate mask
8:    mask $\leftarrow$ 1
9:  else
10:    mask $\leftarrow$ 0
11:  end if
12:  Xmasked $\leftarrow$ mask * (Xrolled -- X)   $\triangleright$ get features
13:  Xfinal $\leftarrow$ $\mathtt{max}$(Xmasked, Xfinal)   $\triangleright$ keep max
14:  $m$leftarrow $m+1$
15:end while
16:$m$leftarrow 0
17:while $mK$ $<$ W do
18:  Xrolled $\leftarrow$ $\mathtt{roll}_{right}$(X, $mK$)
19:  dist $\leftarrow$ $\mathtt{norm}$(X, Xrolled)
20:  if dist $<$ ($\mu$- $\sigma$) then
21:    mask $\leftarrow$ 1
22:  else
23:    mask $\leftarrow$ 0
24:  end if
25:  Xmasked $\leftarrow$ mask * (Xrolled -- X)
26:  Xfinal $\leftarrow$ $\mathtt{max}$(Xmasked, Xfinal)
27:  $m$leftarrow $m+1$
28:end while
29:return $\mathtt{Conv2d}$($\mathtt{Concat}$(X, Xfinal))
Now that we have obtained the estimated mean $\mu$ and standard deviation $\sigma$ of the image, we translate the input feature map $X$ by $m_{K}$ pixels either horizontally or vertically, provided that $m_{K}$ is smaller than the height $H$ and width $W$ of the feature map, as shown in Algorithm 1. This translation operation is used to compare the feature patches that are separated by $N$ steps. In Fig. 5 (b), node $(5,1)$ in the coordinates of $(x,y)$ is compared to nodes $(3,2)$, $(6,2)$, $(7,2)$, $(5,3)$ through ”rolling” to the next node. After the rolling operation, we compute the Euclidean distance between the input $X$ and its rolled version $X_{\text{rolled}}$ to determine whether the two points should be connected. If the distance is less than $\mu-\sigma$, the mask is assigned a value of 1; otherwise, it is assigned a value of 0. This mask is then multiplied by $X_{\textit{rolled}}-X$ to suppress the maximum relative scores between the feature patches that are not considered connected. In Algorithm 1, these values are denoted as $X_{\textit{down}}$ and $X_{\textit{right}}$, respectively. Next, a max operation is performed, and the result is stored in $X_{\textit{final}}$. Finally, after the rolling, masking, and max-relative operations, a final $\mathtt{Conv2d}$ computation is applied.
Through our proposed approach, DGGF constructs a more representative graph structure compared to GGF [37], as it does not connect dissimilar patches (i.e., nodes). Additionally, compared to KNN, DGGF significantly reduces the computational overhead by minimizing adjacency computations during graph construction (whereas KNN must determine the nearest neighbors for each image patch). Furthermore, DGGF does not require the reshaping necessary for performing graph convolution in KNN-based methods [54]. Therefore, DGGF combines the representational flexibility of KNN with the computational efficiency of GGF.
Based on the dynamically constructed graph described above, we connect it to a feedforward neural network with a GeLU activation. Given an input feature $f_{t}^{lc}\in\mathbb{R}^{N\times N}$, the detailed process of updated dynamic grapher is as follows:
$lc=\phi(\mathtt{DynConv}(f_{t}^{lc}))\mathbf{W}_{in})\mathbf{W}_{out}+lr,$ (12)
where $lc\in\mathbb{R}^{N\times N}$ denotes the output graph. $\mathbf{W}_{in}$ and $\mathbf{W}_{out}$ are two weights of the feedforward layers. $\phi$ denotes the GeLU activation.
III-E C3D-RECHead
Conventional 3D object detection models based on point clouds, such as CenterPoint [55], typically regress the bounding box relative to the object’s center. However, in 3D visual grounding based on point cloud sensors, we identify two key limitations of the center-based regression: (1) Higher Point Cloud Density at the Nearest Edge: The nearest object to the autonomous vehicle has a denser point cloud, especially on the side facing the vehicle, which provides a more reliable anchor point for regression, improving localization accuracy. (2) Alignment with Textual Distance Prompts: Many textual prompts in 3D visual grounding include object distances (e.g., ”the car 5 meters ahead”). These distances typically refer to the edge of the object closest to the vehicle, rather than the center. Thus, regressing from this nearest edge point aligns better with semantic and perceptual grounding.
Instead of regressing from the object’s center, we propose C3D-RECHead. Exactly, it performs bounding box regression from the nearest edge point to the vehicle. This adaptation preserves the computational efficiency of CenterPoint while improving localization accuracy in 3D visual grounding tasks. Assume a 3D bounding box $B$ is represented as:
$B=\{p_{c},l,w,h,\theta\},$ (13)
where $p_{c}$ denotes the center of the box, $l,w,h$ denote length, width and height of the box, respectively, and $\theta$ denotes the orientation of the box.
For each bounding box, the eight corner points are:
$P=\{p_{1},p_{2},\dots,p_{8}\},$ (14)
where each $p_{i}$ is computed using the box dimensions and orientation.
Given the sensor position $p_{s}$, the edge set $E$ consists of four candidate edges (assuming a rectangular box base in BEV):
$E=\{e_{1},e_{2},e_{3},e_{4}\},$ (15)
where each edge is defined as two points:
$e_{i}=(p_{i},p_{j}),i\neq j.$ (16)
We select the nearest edge $e^{*}$ by:
$e^{*}=\underset{e_{i}\in E}{\mathtt{argmin}}\ d(e_{i},p_{s}),$ (17)
where $d(e_{i},p_{s})$ is the Euclidean distance from the sensor to each edge.
Figure 6: The difference between CenterPoint-based prediction head and our proposed C3D-RECHead.
III-F Loss Function
Our proposed C3D-RECHead first generates heatmaps to predict the nearest corner location of each detected object, serving as an anchor point for subsequent regression. This approach leverages the increased density of LiDAR and radar point clouds around the closest object, improving localization accuracy. Following the heatmap-based localization, we refine key object properties under the CenterPoint-inspired supervision paradigm. Specifically, our model predicts:
  1. 1.
    Sub-voxel location refinement to enhance localization precision beyond the voxel grid resolution.
  2. 2.
    Height above ground to capture elevation information crucial for differentiating objects in 3D space.
  3. 3.
    3D bounding box dimensions include width, length, and height to characterize object shape and size.
  4. 4.
    Orientation angle to accurately model object rotation in the 3D scene.
Finally, our proposed TPCNet is trained using a multi-task loss function, integrating these regression objectives to optimize performance in 3D visual grounding. The loss formulation is as follows:
${L}_{\rm total}={L}_{\rm hm}+\beta\textstyle\sum\limits_{r\in\boldsymbol{\Lambda}}{L}_{{\rm smooth}-\ell_{1}}(\widehat{\Delta r^{a}},\Delta r^{a}),$ (18)
where ${L}_{\rm hm}$ is the confidence loss supervising the heatmap quality of the center-based detection head using a focal loss [56]; $\boldsymbol{\Lambda}=\left\{x,y,z,l,h,w,\theta\right\}$ indicates the smooth-$\ell_{1}$ loss supervising the regression of the box center (for slight modification based on heatmap peak guidance), dimensions, and orientation; and $\beta$ is the weight to balance the two components of the loss, which is set to 0.25 by default.
IV Experiments and Performance Analysis
Table I: Overall performances on Talk2Radar dataset [37] (Best Performance. Car, Pedestrian and Cyclist provide specialized mAPs. Radar5 denotes the accumulation of five frame radar data.)
Models Sensors Text Encoder Fusion Entire Annotated Area (EAA) Driving Corridor Area (DCA)
Car Pedestrian Cyclist mAP mAOS Car Pedestrian Cyclist mAP mAOS
PointPillars Radar5 ALBERT [57] HDP 18.92 9.79 12.47 13.73 12.91 39.20 10.25 14.93 21.46 20.19
CenterFormer Radar5 ALBERT HDP 17.26 6.79 9.27 11.11 10.79 19.56 9.13 12.03 13.57 13.02
CenterPoint Radar5 ALBERT HDP 18.98 5.30 14.96 13.08 12.20 40.53 8.57 15.66 21.59 20.25
PointPillars Radar5 ALBERT MHCA 5.18 5.76 6.63 5.86 3.58 13.34 4.36 8.79 8.83 7.66
CenterFormer Radar5 ALBERT MHCA 4.53 3.48 4.00 4.00 2.03 8.77 3.52 6.69 6.33 5.92
CenterPoint Radar5 ALBERT MHCA 5.21 4.57 5.13 4.97 3.12 12.70 4.07 7.70 8.16 7.51
MSSG [36] Radar5 GRU [58] - 12.53 5.08 8.47 8.69 7.03 18.93 7.88 9.40 12.07 11.67
AFMNet [59] Radar5 GRU - 11.98 6.87 9.16 9.34 7.72 18.62 8.21 10.06 12.30 11.79
MSSG [36] Radar5 ALBERT - 16.03 5.86 10.57 10.82 8.96 25.79 8.69 12.55 15.68 14.12
AFMNet [59] Radar5 ALBERT - 16.31 6.80 10.35 11.15 9.46 26.82 8.71 12.45 15.99 14.18
EDA [60] Radar5 RoBERTa [50] - 13.23 6.60 8.63 9.49 8.93 23.55 8.80 11.95 14.77 13.07
T-RadarNet [37] Radar5 ALBERT GGF 24.68 9.71 15.74 16.71 14.88 42.58 10.13 17.82 23.51 22.37
TPCNet (ours) Radar5 PointCLIP [48] DGGF 25.92 9.60 16.37 17.30 15.29 45.97 12.73 19.73 26.14 24.73
CenterPoint LiDAR ALBERT HDP 28.16 6.21 17.46 17.28 16.03 43.43 6.87 27.18 25.83 24.93
CenterPoint LiDAR ALBERT MHCA 6.56 5.04 5.33 5.64 4.86 13.60 4.52 7.32 8.48 7.89
MSSG [36] LiDAR GRU - 15.38 7.52 11.67 11.52 9.76 23.27 8.68 13.51 15.15 14.75
AFMNet [59] LiDAR GRU - 16.13 7.68 12.51 12.11 9.92 24.50 9.07 13.87 15.81 15.11
MSSG [36] LiDAR ALBERT - 18.19 7.66 11.63 12.49 10.91 29.61 10.98 14.66 18.42 16.23
AFMNet [59] LiDAR ALBERT - 19.50 7.92 13.56 13.66 12.18 31.68 9.23 18.90 19.94 17.59
EDA [60] LiDAR RoBERTa - 16.10 6.91 12.88 11.96 10.10 25.10 9.28 15.73 16.70 14.91
T-RadarNet [37] LiDAR ALBERT GGF 24.91 12.74 18.67 18.77 17.20 48.98 14.69 27.24 30.30 29.89
TPCNet (ours) LiDAR PointCLIP DGGF 28.93 12.83 20.72 20.83 18.94 49.78 15.25 29.11 31.38 30.14
TPCNet (ours) LiDAR+Radar5 PointCLIP DGGF 32.13 15.99 23.74 23.95 22.01 52.72 17.30 31.66 33.89 31.73
IV-A Settings of Models and Implementations
Model Settings: Besides the PointCLIP [48] that we leverge in our proposed TPCNet, for other models with pre-trained transformer (e.g., ALBERT [57]) for text encoding, we set the token length as 30, uniformly. The number of pillars in the pillar backbone of TPCNet is set as 10 and 32 for radar and LiDAR, respectively. We set the sizes of both LiDAR and Radar Context Agent Feature as $12\times 12$ in default.
For comparison, we firstly select PC-based detectors with various paradigms including PointPillars (pillar-based) [47], CenterPoint (voxel and anchor-free) [55], and CenterFormer (transformer-based) [61], which all fuse the point cloud and textual features between the stage of backbone and FPN, following the same fusion paradigm with our proposed TPCNet. Among the above models, we leverage the SECOND FPN [62] as the multi-scale feature fusion module in PointPillars and CenterPoint. Secondly, the fusion methods of point clouds and text include inductive bias-based HDP [63], attention-based MHCA [64] and graph-based GGF [37] are implemented to compare with our DGGF. Lastly, we also compare proposed TPCNet with dedicated SOTA 3D PC grounding models, including T-RadarNet [37], MSSG [36], AFMNet [59] and EDA [60]. Moreover, various LiDAR and radar fusion approaches, including Multi-Head Cross Attention (MHCA) [52] and Multi-Head Linear Attention (MHLCA) [65], InterRAL [66], L2R Fusion [27] and Adaptive Gated Network (AGN) [24], are compared with our fusion approach as well.
Dataset Settings: We conduct comprehensive training and evaluation of models on the Talk2Radar dataset [37], which encompasses three distinct categories of traffic participants: Car, Cyclist, and Pedestrian, utilizing both 4D radar and LiDAR modalities. Furthermore, to rigorously assess the generalization capabilities of T-RadarNet, we extend our evaluation to the Talk2Car dataset [33]. This dataset, derived from the nuScenes benchmark [67], provides synchronized LiDAR and radar data accompanied by textual prompts referencing object objects, thereby enabling a robust validation of our approach across diverse multi-modal scenarios.
[…]
Models Sensors BEV AP 3D AP
[…]
Baseline [33] L 30.6 24.4 27.9 19.1
MSSG [36] L 27.8 26.1 31.9 20.3
EDA [60] L 37.0 29.8 37.2 20.4
AFMNet [59] L 45.3 33.1 41.9 20.7
T-RadarNet [37] L 52.8 39.9 47.2 30.5
TPCNet (ours) L+R 56.7 44.1 52.3 33.6
Training and Evaluation Settings: For the Talk2Radar dataset, all models are trained using a distributed setup across four RTX A4000 GPUs, with a batch size of 4 per GPU for a total of 80 epochs. The optimization process employs AdamW with an initial learning rate of $1\times 10^{-3}$, modulated by a cosine annealing learning rate scheduler, and a weight decay of $5\times 10^{-4}$. To evaluate 3D visual grounding performance, we utilize two key metrics: Average Precision (AP) and Average Orientation Similarity (AOS), which are computed for both the entire annotated area and the driving corridor. Specifically, we report the mean Average Precision (mAP) and mean Average Orientation Similarity (mAOS) for 3D bounding box localization and orientation estimation, respectively.
For the Talk2Car dataset, we adhere to the baseline configurations outlined in [36]. Specifically, models are trained on four RTX A4000 GPUs with a batch size of 1 per GPU for 20 epochs. Optimization is performed using Stochastic Gradient Descent (SGD) with a momentum of 0.9, a weight decay of $1\times 10^{-4}$, and an initial learning rate of $1\times 10^{-2}$, which is similarly adjusted via a cosine annealing scheduler. The evaluation metric for 3D visual grounding is exclusively based on Average Precision (AP), which quantifies the localization accuracy of the predicted 3D bounding boxes relative to the ground truth annotations.
IV-B Quantitative Results
Overall comparison on Talk2Radar and Talk2Car datasets. As shown in Table I, our proposed TPCNet, which integrates LiDAR and radar, achieves state-of-the-art performance on the Talk2Radar dataset. Compared to unimodal models (LiDAR-only and radar-only), TPCNet demonstrates a significant improvement in both mAP and mAOS metrics. Furthermore, even in its unimodal configuration, TPCNet outperforms the models with similar sensor inputs, highlighting the superiority of our model architecture. In terms of per-class accuracy, the detection performance for Car is slightly higher than that for Cyclist, while Pedestrian exhibits the lowest accuracy. Additionally, on the Talk2Car dataset, TPCNet also surpasses the best-performing LiDAR-only model, demonstrating strong generalization across sensor data distributions.
Table III: Comparison (mAP) of different modalities upon various prompts in the Talk2Radar dataset [37]. R1, R3 and R5 denote the accumulation of one, three and five frames of radar data, respectively. L denotes LiDAR. The experiments of single modality are implemented by T-RadarNet [37] and the performances of multiple modalities are implemented by TPCNet.
Prompt Motion Depth Velocity
Sensors Car Pedes Cyclist Car Pedes Cyclist Car Pedes Cyclist
R1 26.1 5.3 11.1 32.7 12.7 24.7 26.9 14.0 18.5
R3 35.9 11.1 15.4 40.7 18.7 31.7 33.8 20.5 25.8
R5 36.7 11.5 16.3 42.5 19.6 32.0 35.5 20.8 25.7
L 33.6 8.8 13.6 45.7 20.5 38.6 12.6 7.6 8.2
R1+L 37.6 13.4 18.9 48.8 22.5 41.9 38.7 26.5 29.8
R3+L 37.7 12.8 19.7 48.2 24.1 42.5 39.6 27.8 30.9
R5+L 38.9 14.1 21.2 49.7 25.7 42.7 40.8 28.6 32.0
[…]
Methods Talk2Radar Talk2Car
[…]
Bi-MHCA [52] 29.14 25.87 49.7 33.2
Bi-MHLCA [65] 24.53 23.92 48.1 30.9
InterRAL [66] 26.45 24.23 50.8 31.3
L2R Fusion [27] 25.26 22.94 48.5 30.1
AGN [24] 27.84 25.03 50.1 31.0
BACA (ours) 28.92 26.07 52.3 33.6
Table V: Comparison of efficiency on fusion methods of LiDAR and radar features with different sizes. $N_{h}$ denotes the number of heads in cross attention, which is only for the attention-based fusion method. FPS is the frame per second of TPCNet equipped with various fusion methods on single RTX A4000 GPU.
Methods Params $\downarrow$ FLOPs $\downarrow$ FPS $\uparrow$
Stage 3 of pseudo image: $D$=64, $H$=80, $W$=80, $N_{h}$=4
Stage 4 of pseudo image: $D$=128, $H$=40, $W$=40, $N_{h}$=8
Stage 5 of pseudo image: $D$=256, $H$=20, $W$=20, $N_{h}$=8
Non-Attention-based Fusion Methods
AGN [24] 147.58K 1.89G
590.08K 1.89G 19.7
2.36M 1.89G
Attention-based Fusion Methods
Bi-MHCA [52] 0.33M 10.70G
0.13M 1520.44M 4.8
0.52M 373.56M
Bi-MHLCA [65] 131.14K 1.68G
262.27K 838.86M 9.5
524.54K 419.43M
BACA (ours) 0.64K 46.61M
Agent Feature Len 1.28K 23.33M 18.2
=8$\times$8 2.56K 6.81M
BACA (ours) 0.64K 164.62M
Agent Feature Len 1.28K 82.40M 16.2
=16$\times$16 2.56K 21.75M
BACA (ours) 0.64K 206.41M
Agent Feature Len 1.28K 103.32M 13.0
=18$\times$18 2.56K 27.04M
Performances of TPNet based on two input sensors for different types of prompts. Table III presents the performance of TPCNet with different sensor inputs when various types of prompts are processed. Overall, we observe that for text prompts containing motion or velocity information, radar generally outperforms LiDAR. In contrast, for prompts involving object depth, the LiDAR-only model performs better than the radar-only model. However, the fusion of radar and LiDAR significantly outperforms both unimodal approaches. Additionally, aggregating five frames of radar point clouds yields better performance than using three or one frame. This demonstrates the strong complementarity between radar and LiDAR.
Table VI: Comparison of performances between C3D-RECHead and CenterPoint-based prediction head.
Methods Entire Annotated Area Driving Corridor Area
 mAP  mAOS  mAP  mAOS
CenterPoint [55] 22.70 20.55 31.92 30.99
C3D-RECHead (ours) 23.95 22.01 33.89 31.73
Performances of various fusion methods of LiDAR and radar under the settings of TPCNet. As TABLE IV shows, under the TPCNet architecture, our proposed BACA outperforms other LiDAR-radar fusion methods. Overall, it can be observed that the global fusion methods based on cross-attention achieve superior performance when compared with convolution-based fusion approaches. Specifically, the Bi-MHCA-based fusion method achieves the highest mAP, followed closely by our proposed BACA, which performs slightly worse than Bi-MHCA but surpasses Bi-MHLCA. Moreover, on the Talk2Car dataset, our BACA achieves the best performance.
Table VII: Statistics of mAP for predicted objects by TPCNet with the respect with depth upon 5-frame radar, LiDAR and the combination of radar and LiDAR.
Sensors Objects 0-10 (m) 10-20 20-30 30-40 40-50 50+
Radar5 Car 43.62 45.19 25.32 20.13 11.28 2.20
Pedestrian 16.79 8.11 3.98 3.09 4.89 0.0
Cyclist 31.55 14.23 16.17 5.74 3.23 1.22
LiDAR Car 46.78 49.50 31.72 26.45 14.19 8.12
Pedestrian 19.55 9.79 7.88 6.57 5.48 0.0
Cyclist 36.92 17.63 20.11 7.53 6.11 3.74
LiDAR Car 51.67 54.69 36.55 29.87 18.56 10.59
+ Pedestrian 21.41 11.16 7.92 7.10 5.79 0.0
Radar5 Cyclist 38.55 21.27 22.79 10.57 6.20 4.12
The comparison of efficiency between various fusion methods for LiDAR and radar. As shown in Table IX, we have compared the complexity and inference speeds of non-attention-based fusion methods (convolution-based AGN) and attention-based fusion methods for processing multi-scale feature maps from the two modalities. We observe that our proposed BACA has significantly fewer parameters than other fusion methods, with a 2 to 3 orders of magnitude reduction in the parameter count when compared to the other three methods. Moreover, the FLOPs of BACA are substantially lower than those of Bi-MHLCA, Bi-MHCA, and AGN. At the same time, BACA achieves a much higher FPS than other attention-based fusion methods and is very close to the convolution-based AGN in terms of inference speed. Furthermore, when setting the agent feature map size to 8×8, 16×16, and 18×18, the increase in FLOPs remains relatively small, demonstrating the strong scalability of BACA.
Table VIII: Ablation experiments of TPCNet with various settings on Talk2Radar dataset.
Methods EAA DCA
Car Ped Cyc Car Ped Cyc
Dynamic Gated Graph Fusion
Dynamic $\rightarrow$ Static [53] 29.97 13.98 21.40 50.79 15.89 28.63
MaxPool $\rightarrow$ AvgPool 30.19 14.67 20.55 51.41 16.36 29.74
Dynamic + MaxPool 32.13 15.99 23.74 52.72 17.30 31.66
Text Encoders
Bi-GRU [58] 25.97 11.80 17.11 46.59 12.58 24.77
RoBERTa [50] 30.03 13.79 21.14 51.88 15.13 28.10
ALBERT [57] 29.98 14.01 20.97 50.29 16.42 29.77
PointCLIP [48] 32.13 15.99 23.74 52.72 17.30 31.66
Fusion Methods of Text and Point Cloud
HDP [63] 22.57 14.68 16.39 43.78 10.57 23.66
MHCA [64] 11.23 6.54 8.96 18.20 7.96 10.29
MHLCA [65] 9.12 6.74 7.05 16.46 8.41 10.10
GGF [37] 28.79 13.10 19.87 48.92 14.22 28.67
DGGF 32.13 15.99 23.74 52.72 17.30 31.66
The comparison between CenterPoint-based prediction head and our proposed C3D-RECHead. As shown in Table VI, we comprehensively evaluate the accuracy of 3D object detection and orientation angle estimation using two different detection head paradigms on the Talk2Radar dataset. Our findings indicate that the proposed C3D-RECHead exhibits a significant performance advantage over CenterPoint, achieving notable improvements in both mAP and mAOS.
[…]
Prompt Motion Depth Velocity
Methods Car Pedes Cyclist Car Pedes Cyclist Car Pedes Cyclist
[…]
BACA 38.9 14.1 21.2 49.7 25.7 42.7 40.8 28.6 32.0
The comparison on the architecture of models. As Fig. 9 shows, we have compared T-RadarNet and TPCNet, with the models taking radar and LiDAR as single-modal inputs, respectively, to analyze the effectiveness of the model architecture. Our results show that, regardless of whether radar or LiDAR is used as the input, TPCNet consistently outperforms T-RadarNet in localization accuracy across all object categories and at various distance intervals from the ego vehicle. This demonstrates the architectural superiority of TPCNet over T-RadarNet.
Figure 7: Comparison between TPCNet and T-RadarNet by depth based on the input of uni-modality.
Statistics of object prediction accuracy at different depths. We partition the predicted objects into six depth intervals, each spanning 10 meters relative to the ego vehicle. As shown in Table VII, the fusion of LiDAR and radar significantly enhances object detection accuracy, particularly for distant objects. The fused model outperforms both LiDAR-only and radar-only modalities, demonstrating the necessity of multi-modal fusion.
IV-C Ablation Experiments
Ablation studies of vital modules in TPCNet. As shown in TABLE VIII: (1) We first conduct an ablation study on the individual components of the DGGF module. Our analysis revealed that dynamic graphs achieve higher accuracy in scene object recognition when compared with static graphs, indicating that dynamic graphs handle environmental redundancy more effectively. Furthermore, in constructing edges between graph nodes, MaxPool outperforms AvgPool in capturing salient inter-node relationships.
(2) We compare the performance of different text encoders on the 3D visual grounding task. Transformer-based encoders significantly outperform Bi-GRU, demonstrating the efficiency of the attention mechanisms in extracting meaningful textual information. However, PointCLIP, which leverages contrastive learning between point clouds and text, outperforms RoBERTa and ALBERT, highlighting its advantage in embedding textual information with 3D point cloud-conditioned features, making it particularly suitable for 3D point cloud-based visual grounding tasks.
(3) We replace different text-point cloud feature fusion modules in TPCNet. Our proposed DGGF outperforms both GGF and HDP, which rely on point-to-point fusion. Additionally, we observe that two global cross-attention-based modules, MHCA and MHLCA, perform poorly. This was primarily due to the excessive noise in the point cloud features (especially from radar), causing the model to overemphasize non-object regions and resulting in a high false positive rate.
Figure 8: Visualization of predicted results by our proposed TPCNet in Talk2Radar dataset.The first row presents the prediction results from the view of camera while the second row shows the results from Bird’s Eye View (BEV).
Comparison of performances on cross attention between LiDAR and radar. TABLE IX shows the results of ablation studies in the BACA module. We compared two model configurations: one where radar features were used as the query while LiDAR features served as the key and value, and the other where LiDAR was used as the query while radar was not incorporated as the key or value. Our findings indicate that when radar served as the primary information source, supplemented by the fine-grained 3D features of LiDAR, TPCNet achieves superior performance in the prompts containing motion and velocity information. However, for samples with depth-related prompts, the fusion method that uses LiDAR as the query outperforms the radar-query configuration. Among the three approaches, BACA demonstrates the best performance across all prompt types. This result highlights the effectiveness of our proposed BACA module in leveraging bidirectional fusion to enrich the environmental context features required by each modality, thereby enhancing its ability to accurately identify the object corresponding to the textual prompt.
Figure 9: Prediction results by LiDAR-only, radar-only and the fusion of LiDAR and radar based on the TPCNet model.
Figure 10: Comparison of various models from the views of RGB camera.
Figure 11: Heatmaps of our proposed C3D-RECHead (first row) and CenterPoint head (second row). Red bounding boxes denote the predicted objects while yellow bounding boxes denote ground truth.
IV-D Visualization and Discussion
Fig. 8 illustrates the prediction results of TPCNet on the Talk2Radar dataset. As shown, TPCNet effectively captures the objects of varying sizes corresponding to textual prompts across different scenarios, even in complex point cloud environments with clutter. The model demonstrates robust performance in detecting both single and multiple objects, as well as objects at varying distances, including those in close proximity and farther away.
Fig. 9 shows the performance of TPCNet under three settings: LiDAR-only, radar-only, and fusion of both sensors. For the first sample, we observe that the radar-based prediction produces false positive detections, while the LiDAR-based model yields accurate predictions but lacks the level of refinement achieved by sensor fusion. In the second-row sample, our proposed fusion method consistently outperforms the two single-sensor approaches. For the final sample, both the LiDAR-only and radar-only models exhibit varying degrees of false positives due to clutter, whereas the fusion-based approach effectively suppresses false positives and accurately localizes the objects.
Fig. 10 compares the performance of three models: T-RadarNet, which relies solely on LiDAR input; TPCNet-AGN, which integrates LiDAR and radar features using the AGN module; and TPCNet-BACA, which employs BACA for dual-sensor feature fusion. In the first sample, where the textual prompt refers to four objects, T-RadarNet is only able to detect the two closest objects while failing to localize the two pedestrians further ahead. TPCNet-AGN successfully identifies all four objects but fails to accurately localize one of them. In contrast, our proposed TPCNet-BACA correctly detects and localizes all four prompted objects. For the second sample, where the prompt contains both distance and orientation information, T-RadarNet accurately localizes the correct object. However, TPCNet-AGN mistakenly identifies the truck in front as the object and fails to correctly localize the car. In the third sample, T-RadarNet incorrectly identifies a car on the front right as the referenced object, while TPCNet-AGN produces an additional false positive detection. Overall, our proposed TPCNet-BACA demonstrates superior accuracy in object identification and localization.
As Fig.11 shows, we visualize the heatmap of two prediction heads, including the CenterPoint-based prediction head and our proposed C3D-RECHead. We find that our proposed C3D-RECHead successfully focuses on the corner point closest to the ego vehicle, which leads to more precise bounding box predictions compared to the CenterPoint head. The predicted bounding boxes from C3D-RECHead align more closely with the ground truth, demonstrating that performing 3D visual grounding based on the nearest edge yields more stable results than anchoring to the object’s center.
V Conclusions
This paper focuses on 3D visual grounding for autonomous driving and embodied perception. We pioneeringly propose a novel model, TPCNet, which enables 3D detection of referential objects in a scene by interpreting textual prompts to guide the dual-sensor fusion of LiDAR and radar. In TPCNet, we first introduce the Bidirectional Agent Cross Attention (BACA) module, which efficiently fuses LiDAR and radar features at the minimal cost while preserving the dynamic characteristic paradigm of the attention mechanism, allowing for low-cost, dynamic fusion of sensor features with textual prompt features. Secondly, we propose a dynamic graph-based fusion approach for point clouds and text features, termed Dynamic Gated Graph Fusion (DGGF). DGGF dynamically extracts point cloud regions of interest based on text embeddings and outperforms other fusion methods. Finally, to enhance 3D visual grounding performance and align with the prompts containing object depth information, we propose C3D-RECHead, which achieves superior performance compared to CenterPoint. Overall, TPCNet achieves state-of-the-art 3D visual grounding performance on both the Talk2Radar and Talk2Car datasets, outperforming existing models.